20 men or 30 boys can do a certain job in 20 days. how long will it take 6 men and 15 boys to do the same job ?
Answer by ankor@dixie-net.com(22740)

20 men or 30 boys can do a certain job in 20 days.
Find how many man-days or boy-days to do the job
20*20 = 400 man-days
or
20*30 = 600 boy-days
therefore
400md = 600 bd
m = 600bd/400d
m = 1.5b
a man accomplishes 1.5 times as much as a boy in 1 hr
:
how long will it take 6 men and 15 boys to do the same job ?
let t = no. of days to do this (in terms of boy-days)
6t(1.5) + 15t = 600
9t + 15t = 600
24t = 600
t = 600/24
t = 25 days to complete the job